# Multi-label Scene Classification for Autonomous Vehicles: Acquiring and Accumulating Knowledge from Diverse Datasets

### v2

@@ V Implementation Details
The proposed KAA-CAL learning system was implemented using PyTorch 1.10.0 on a server equipped with an Nvidia Tesla V100 featuring 32 GB of memory. Model training utilized the AdamW optimizer, a batch size of 32, and a maximum of 100 training epochs. Input images were resized to $224\times 244$ and augmented with random rotations, crops, and pixel normalization.
V-A KAA Implementation
The deep encoder for KAA in Eq. (5) was initialized by loading the weights pretrained on ImageNet-22k [32]. The projectors in Eq. (6) and the classifiers in Eq. (7) were initialized using the default Xavier initialization method [33]. Then, learnable model parameters were progressively optimized using the method delineated in Algorithm 1.
The learning rate for KAA implementation was adjusted in a two-phase approach. It began with a 10-epoch warm-up, where the learning rate linearly increased from an initial value of 1e-6 to 5e-4. Following this, the learning rate decayed to a final value of 1e-5 using a cosine schedule. Upon convergence at the end of all training epochs, the model from the final epoch was chosen as the ultimate model.
V-B CAL Implementation
In implementing CAL, a per-iteration budget ($b_{j}$) of 200 images (approximately 1% of the single-label training data in DSI) was allocated for annotating the data selected using Algorithm 2. This process ran for up to five iterations to evaluate the efficiency of CAL.
The data annotation process confronted a challenge. Not all images can be reliably annotated with all the seven driving scene attributes. For example, the weather condition in some night-time images was hard to identify, and annotators were not always confident about the road function for some scene images. When a reliable label was unavailable, a class label of “-1” was assigned. Critically, these labels were excluded from the loss function evaluation to ensure training integrity.
Aiming to retain knowledge learned from KAA, the first three stages of the deep encoder were frozen for finetuning multitask classification. The learning rate scheduler was similarly configured for the CAL implementation: it began with a 10-epoch warm-up, with the learning rate linearly increasing from 1e-6 to 1e-3. After that, the learning rate decayed to a final value of 1e-5 based on a cosine schedule.
@@ VI Experiments and Results
This section presents experimental studies, aiming to verify the merits of the proposed KAA-CAL learning system. First, Subsections A-C will assess the advantages of KAA’s designs and the performance of the resulting foundation model. Then, Subsection D will evaluate the efficiency and effectiveness of CAL that adapts the foundation model for multitasking. Upon those analyses, an ablation study will be further performed in Subsection E to quantify the respective and joint contributions of learning and adaptation, supplemented by examples presented in Subsection F. In the end, Subsection G will compare the multi-label classification model attained via KAA-CAL to its variants and state-of-the-art (SOTA) multitask models on public datasets.
@@ VI-A Deep Encoder Selection
Vision transformers (ViT) demonstrate unique advantages over convolutional neural networks (CNNs) in learning deep features for image classification tasks. In selecting a deep encoder for the KAA-CAL learning system, the study compared SOTA transformer-based models, including Swin-S, Swin-B, ViT-B, and ViT-L. This comparison spanned the seven independent image classification tasks, evaluating models from the perspectives of both classification accuracy and model complexity. The results are summarized in TABLE I, with the highest accuracy achieved on each task being bolded.
@@
As expected, Swin-B outperforms Swin-S on all seven tasks, and the same pattern holds for ViT-L compared to ViT-B. This confirms that the more complex versions of the models (Swin-B and ViT-L) achieve better performance than their less complex counterparts. While both are powerful vision transformers, Swin transformers address several limitations of the original ViT, exhibiting improved computational efficiency and the capability for multi-scale feature learning. TABLE I shows that Swin-B outperformed ViT-L on five out of seven tasks with a margin ranging from 0.1% to 4.9%. Importantly, the FLOPs for Swin-B are only 20% of ViT-L. Given the highest average accuracy and more affordable model complexity, Swin-B is chosen as the deep encoder architecture for the classification model.
@@ VI-B Effectiveness of Stage-wise Knowledge Distillation
Sections IV-A3 and IV-A4 explained that, when KAqN continues acquiring new knowledge, KAA retains the already acquired knowledge accumulated in KAcN through stage-wise feature-based KD, and the intensity of this regularization is moderated by PARI. To demonstrate the effectiveness of the knowledge retention design for KAA, an ablation study about KD was performed with results summarized in TABLE II.
@@
The Baseline model in TABLE II is learned solely by a student network because both stage-wise KD and PARI are removed. It results in an average accuracy of 92.7%, representing a 2.5% decrease compared to our model (Model IV). It is suspected that the gap will increase when the KAA attempts to acquire more diverse knowledge from many datasets.
Model I adds last-stage KD to the baseline model, analogous to a teacher who provides less guidance to the student’s learning. This model achieves a negligible performance gain of merely 0.1%. Conversely, Model III incorporates stage-wise KD into the baseline model, leading to a substantial average accuracy increase of 1.8%. This comparison highlights the importance of stage-wise KD, especially since features of different scales contain comprehensive information about driving scenes.
Model II augments last-stage KD with PARI, which improves average accuracy by 1.2% over Model I. Similarly, Model IV pairs PARI with stage-wise KD, increasing Model III’s well-performed average accuracy by an additional 0.7%. These observations verify the effectiveness of dynamically adjusting the weight for new knowledge acquisition according to specific learning needs, regardless of the KD method.
@@ VI-C Performance of the Foundation Model
After verifying the strengths of KAA’s architecture design, the experimental study continued to assess the resulting foundation model’s performance. Fig. 4 compares the foundation model to the seven individual monotask models. As illustrated, the foundation model is comparable to the individual monotask models in task accuracy, with variations ranging from -3.9% to 3.2% across the seven tasks. Specifically, the foundation model is 3.2% more accurate than the monotask model for identifying Grade-separated Infrastructure. However, the foundation model’s accuracy in identifying Work Zone is 3.9% lower. This drop can be attributed to the challenge of image-level classification in capturing the fine-grained features critical for the identification of scenes within work zones, such as safety cones and work zone signs. In contrast, approaches based on object detection or semantic segmentation [34, 35] are generally more capable of detecting such localized features. For the remaining five tasks, their performance differences are within ±$\pm 1$%. The foundation model can substitute for individual models dedicated to different classification tasks, as its comparable performance demonstrates.
Figure 4: Comparison of the foundation model with the seven monotask models in terms of classification accuracy
@@ VI-D Effectiveness of Consistency-based Active Learning
As introduced in Section IV-B, CAL is designed to adapt the foundation model to multitask learning, addressing the limitations inherent in KAA’s monotask learning approach. To demonstrate its effectiveness, CAL was compared to its variants and SOTA deep active learning methods, detailed as follows.
@@
  • Ran-I. In this approach, the deep encoder for the classification model is initialized with the weights pretrained on ImageNet-22k. Samples are randomly selected from each class within the unlabeled pool $\mathcal{D}_{U}$ and subsequently annotated for training the model. Compared to CAL, Ran-I does not leverage the driving scene knowledge offered by the foundation model, nor does it employ the feature consistency-based criterion for sampling.
@@
  • Ran-D. This approach uses the same sampling method as Ran-I, but it adopts the foundation model as its deep encoder. During training, the encoder’s first three stages were frozen.
@@
  • Core-Set-I [24]. This is a diversity-based active learning approach proposed in [24]. It’s deep encoder is initialized with the weights pretrained on ImageNet. In the first iteration, the model is refined with a starting training set of $b_{1}$ samples taken from the unlabeled data pool ($\mathcal{D}^{U}$), helping expose the encoder to some multi-label training data. After that, in each subsequent iteration, a k-Center-Greedy strategy selects additional diverse samples that represent various clusters in the unlabeled pool.
@@
  • LLAL-I [23]. This active learning approach follows the same initialization as Core-Set-I, but it introduces an additional loss prediction component developed in [23]. This component is learnable, and it estimates a pseudo-loss value for each sample, aiming to approximate the true task loss. During sample selection from the unlabeled pool, the module prioritizes samples with the highest pseudo-loss values, under the assumption that they represent the most informative samples to learn from.
@@
Fig. 5 compares the proposed CAL method with those introduced above, presenting both the task-level accuracy (a-g) and the average accuracy (h). The plots illustrate test accuracy values before adaptation (the training set $\mathcal{D}^{L}$ has 0% labeled data from $\mathcal{D}^{U}$) and over five iterations of adaptation ($\mathcal{D}^{L}$ consists of 1% to 5% of data sampled from $\mathcal{D}^{U}$). Prior to adaptation for multitask learning, models initialized with the foundation model’s weights (i.e., CAL and Ran-D) demonstrate a higher average accuracy than those initialized with weights pretrained on ImageNet (i.e., Ran-I, LLAL-I, and Core-set-I), with a margin of 31.3%. The performance gap between these two groups before model adaptation highlights the benefit of utilizing the foundation model for driving scene identification, as it possesses specialized knowledge relevant to this application.
@@
After the first iteration, while all methods improve their task performance, CAL is the top one. It surpasses the best competing method’s accuracy by a significant margin of 3.3% to 10.0% across all tasks, except for the task of identifying Time of Day (Fig. 5(a)). Effectively, CAL brings the classification model’s average accuracy up to 86.2% after just one iteration of adaptation. With another iteration, CAL boosts the average accuracy to 90.1%, with task accuracy ranging from 83.6% to 96.3%. Through additional iterations, CAL further enhances the accuracy for identifying Weather and Road Function, thus increasing the average accuracy to 92.7%. Figure 5(h) reveals CAL’s superior performance and cost-effectiveness. It consistently outperforms the best competing method in every iteration, with margins ranging from 5.0% to 12.9%. Importantly, CAL has achieved its major improvements in just two iterations, while all other methods require more iterations and so more training data for adaptation.
@@
Ran-D and CAL differ in their sampling strategies. Ran-D uses a simple stratified random sampling strategy, whereas CAL samples data based on features consistency measurement between unlabeled data and the test set. CAL consistently improves task accuracy over iterations, whereas Ran-D lowers the accuracy in identifying Road Function and Grade-separated Infrastructure. Moreover, CAL achieves a higher saturated accuracy than Ran-D on all tasks. Consequently, CAL clearly dominates Ran-D in Fig. 5(h), demonstrating the importance of CAL’s selective sampling strategy.
@@
Ran-I, LLAL-I, and Core-Set-I do not benefit from the knowledge attained by the foundation model. Therefore, they need more iterations and so more annotation cost than CAL to improve task performance. Among those methods, Core-Set-I seems the most competitive. After five iterations, Core-Set-I reached task accuracy comparable to CAL’s in identifying Weather-related Road Condition and Grade-separated Infrastructure, but CAL still beat Core-Set-I on the remaining five tasks by a margin ranging from 1.3% to 13.9%. Consequently, Core-Set-I’s average accuracy after five iterations is still 5% lower than that of CAL. CAL’s superior efficiency and consistently good performance across all tasks, as compared to Core-Set-I, highlight the importance of leveraging the foundation model’s knowledge obtained through KAA.
@@
Foundation 73.9 56.8$\uparrow$ 65.8 52.6$\uparrow$ 73.0 60.3$\uparrow$ 55.3  2.2$\uparrow$ 48.1  6.0$\uparrow$ 66.4 13.1$\uparrow$ 92.8 28.2$\uparrow$ 67.9 31.3$\uparrow$
with Adaptation ✓(2 iters) 90.0 16.1$\uparrow$ 84.1 18.3$\uparrow$ 83.6 10.6$\uparrow$ 96.3 41.0$\uparrow$ 86.9 38.8$\uparrow$ 93.8 27.4$\uparrow$ 96.3 3.5$\uparrow$ 90.1 22.2$\uparrow$
with Adaptation ✓(5 iters) 92.12.1$\uparrow$ 90.96.8$\uparrow$ 88.65.0$\uparrow$ 97.10.8$\uparrow$ 88.21.3$\uparrow$ 94.20.4$\uparrow$ 97.51.2$\uparrow$ 92.72.6$\uparrow$
@@
Observations from Fig. 5 also indicate that the CAL method significantly benefits from the foundation model. The foundation model not only sets up a higher starting point for adaptation but also allows CAL to utilize its superior feature extraction ability to search and identify new training data from the unlabeled pool for ongoing learning and refinement. That is, a suitable active learning method, combined with a powerful foundation model, is the underlying reason for the efficiency and effectiveness of model adaptation.
@@ VI-E Ablation Study
KAA and CAL have now been verified as advantageous designs for a learning system. KAA builds a foundation model for driving scene identification by acquiring and accumulating knowledge from different single-label datasets via monotask learning. The CAL module then adapts this foundation model for multitasking. An ablation study was conducted to gauge their respective contributions, with results summarized in TABLE IV.
The baseline model in this experimentation is a multitask classification model whose deep encoder is the Swin-B pretrained on ImageNet. This baseline model achieves task accuracies ranging from 17.1% to 64.4%, resulting in an average accuracy of 36.6%. The low accuracy observed indicates a lack of knowledge regarding driving scene identification.
Adoption of the foundation model as the deep encoder boosts the average accuracy to 67.9%, representing a substantial increase of 31.3%. This gain is attributed to the comprehensive knowledge regarding driving scenes acquired by KAA from diverse single-label datasets. For instance, the accuracy in identifying Road Function, a unique attribute specific to driving scenes, had an increase of 60.3%. Furthermore, the task accuracies for identifying the environmental attributes, Time of Day and Weather, each rise by over 50%. Consequently, the task accuracies now span from 48.1% to 92.8%. However, there is still clear room for improvement. The gains in identifying Weather-related Road Condition and Intersection are limited, primarily due to the shift in data distribution in real-world scenarios compared to the marginal distributions in the single-labeled datasets.
CAL, designed to address the limitation of KAA’s monotask learning approach, effectively improves the average accuracy to 90.1% with only two iterations, corresponding to another increase of 22.2%. At the task level, accuracies have reached the range from 83.6% to 96.3%. With three additional iterations, the task accuracies in identifying Weather and Road Function are improved by at least 5%, further boosting the average accuracy to 92.7%. The efficient improvement in the accuracy of multitask-based driving scene classification demonstrates CAL’s effectiveness in integrating knowledge acquired through monotask learning on single-label datasets into knowledge suitable for multi-label scene identification.

### v3

@@ V Implementation Details
The proposed KAA-CAL learning method was implemented using PyTorch 1.10.0 on a server equipped with an Nvidia Tesla V100 featuring 32 GB of memory. Model training utilized the AdamW optimizer, a batch size of 32, and a maximum of 100 training epochs. Input images were resized to $224\times 244$ and augmented with random rotations, crops, and pixel normalization.
The deep encoder for KAA in Eq. (4) was initialized with the weights pretrained on ImageNet-22k [30]. The projectors in Eq. (5) and the classifiers in Eq. (6) were initialized using the default Xavier initialization method. Then, learnable parameters were progressively optimized using the method delineated in Algorithm 1. The learning rate for KAA implementation was adjusted in two phases. The first phase is a 10-epoch warm-up, where the learning rate linearly increased from an initial value of 1e-6 to 5e-4. In the second phase, the learning rate decayed to a final value of 1e-5 using a cosine schedule. Upon convergence at the end of all training epochs, the model from the final epoch is the foundation model.
In implementing CAL, a per-iteration budget ($b_{j}$) of 200 images, approximately 1% of the single-label training data in DSI, was allocated for annotating the data recommended by Algorithm 2. This process ran for up to five iterations to evaluate the efficiency of CAL. The data annotation process confronted a challenge - not all images can be reliably annotated with all the seven driving scene attributes. For example, annotators found the weather condition in some nighttime images was hard to identify. When a reliable label was unavailable, a sentinel value “-1” was assigned. Such labels were excluded from the loss function evaluation to ensure training integrity. Aiming to retain knowledge learned from KAA, the first three stages of the deep encoder were frozen for finetuning multitask classification. The learning rate scheduler for the CAL implementation was similarly configured.
@@ VI Experiments and Results
Experimental studies were conducted, aiming to verify the merits of the proposed KAA-CAL learning method.
@@ VI-A Deep Encoder Selection
Various deep network architectures are available to choose from for the deep encoder. Vision transformers (ViT) demonstrate unique advantages over convolutional neural networks in learning deep features for image classification. This study compared SOTA transformer-based models, including Swin-S, Swin-B, ViT-B, and ViT-L. The comparison spanned the seven independent image classification tasks, evaluating models from the perspectives of both classification accuracy and model complexity. The results are summarized in TABLE I.
@@
The top performance for each task is bolded.
As expected, Swin-B outperforms Swin-S on all seven tasks, and the same pattern holds for ViT-L compared to ViT-B. This confirms that the more complex transformers (Swin-B and ViT-L) achieve better performance than their less complex counterparts. While both are powerful, Swin transformers address limitations of the original ViT, exhibiting improved computational efficiency and the capability for multi-scale feature learning. TABLE I shows that Swin-B outperformed ViT-L on five out of seven tasks with a margin ranging from 0.1% to 4.9%. Importantly, the FLOPs for Swin-B are only 20% of ViT-L. Given the highest average accuracy and more affordable model complexity, Swin-B is chosen as the deep encoder architecture for the classification model.
@@ VI-B Effectiveness of Stage-wise Knowledge Distillation
The ability to retain previously acquired knowledge while learning new information is essential for the multi-attribute scene identification model to be scalable. Sections IV-A4 and IV-A5 explained that, when KAqN continues acquiring new knowledge, the already acquired knowledge accumulated in KAcN is retained through stage-wise feature-based KD, and the intensity of this regularization is moderated by PARI. To demonstrate the effectiveness of this knowledge retention design for KAA, an ablation study about KD was performed with results summarized in TABLE II.
@@
The Baseline model in TABLE II is learned solely by a student network because both stage-wise KD and PARI are removed. It results in an average accuracy of 92.7%, representing a 2.5% decrease compared to our model (Model IV). This gap will increase when the KAA attempts to acquire more diverse knowledge from many datasets.
Model I adds the last-stage KD to the baseline model, analogous to a teacher who provides less guidance to the student’s learning. This model achieves a negligible performance gain of merely 0.1%. Conversely, Model III incorporates stage-wise KD into the baseline model, leading to an average accuracy increase of 1.8%. This comparison highlights the importance of stage-wise KD, since features of different scales offer comprehensive information about driving scenes.
Model II augments the last-stage KD with PARI, which improves average accuracy by 1.2% over Model I. Similarly, Model IV pairs PARI with stage-wise KD, increasing Model III’s well-performed average accuracy by an additional 0.7%. These observations verify the importance of dynamically adjusting the weight for new knowledge acquisition according to specific learning needs, regardless of the KD method.
@@ VI-C Performance of the Foundation Model
After verifying the strengths of KAA’s learning architecture design, the experimental study continued to assess the resulting foundation model’s performance. Fig. 5 compares the foundation model to the seven individual monotask models. As illustrated, the foundation model is comparable to the individual monotask models in task accuracy, with variations ranging from -3.9% to 3.2% across the seven tasks. Specifically, the foundation model is 3.2% more accurate than the monotask model for identifying Grade-separated Infrastructure. However, the foundation model’s accuracy in identifying Work Zone is 3.9% lower. This drop can be attributed to the challenge of image-level classification in capturing the fine-grained features critical for identifying work zones, such as safety cones and work zone signs. Approaches based on object detection or semantic segmentation (e.g., [31]) are generally more capable of detecting such localized features. For the remaining five tasks, their performance differences are within $\pm 1$%. The foundation model can substitute for individual models dedicated to different classification tasks, as its comparable performance demonstrates.
Figure 5: The foundation model vs. corresponding monotask models
@@ VI-D Effectiveness of Consistency-based Active Learning
The foundation model, obtained from diverse single-label datasets via monotask learning, must quickly adapt to downstream tasks. CAL proposed in Section IV-B is designed for this purpose, addressing the limitations inherent in KAA’s monotask learning approach. To demonstrate its effectiveness, CAL was compared to its variants and SOTA deep active learning methods, detailed as follows.
@@
  • Ran-I’s deep encoder is initialized with the weights pretrained on ImageNet-22k. Samples for each class are randomly selected from the unlabeled pool and then annotated. Compared to CAL, Ran-I does not leverage the driving scene knowledge offered by the foundation model, nor does it employ the feature consistency-based criterion for sampling.
@@
  • Ran-D uses the same sampling method as Ran-I, but it adopts the foundation model as its deep encoder.
@@
  • Core-Set-I is a diversity-based active learning approach proposed in [23]. Its deep encoder is initialized with the weights pretrained on ImageNet-22k. In the first iteration, the model is refined with an initial training set, helping expose the encoder to some multi-label training data. In each subsequent iteration, a k-Center-Greedy strategy selects additional samples to represent various clusters in the unlabeled pool.
@@
  • LLAL-I follows the same initialization as Core-Set-I, but it introduces an additional loss prediction component [22], which learns to estimate a pseudo-loss value for each sample. In selecting samples from the unlabeled pool, those with the highest pseudo-loss values are prioritized, as they represents the most informative samples to learn from.
@@
Fig. 6 compares the proposed CAL method with those introduced above, presenting the task-level accuracies in plots a-g and the average accuracy in plot h. Each plot illustrates the test accuracy values before adaptation (the training set $\mathcal{D}^{L}$ has 0% labeled data from $\mathcal{D}^{U}$) and over five iterations of adaptation. Each iteration selects 1% of data from the unlabeled pool. Fig. 6(h) shows that, prior to adaptation, models that are initialized with the foundation model’s weights (i.e., CAL and Ran-D) demonstrate a higher average accuracy than those initialized with weights pretrained on ImageNet (i.e., Ran-I, LLAL-I, and Core-set-I), with a margin of 31.3%. The initial performance gap between these two groups highlights the benefit of utilizing the foundation model for driving scene identification, as it possesses the knowledge about the seven classification tasks.
@@
After the first iteration, CAL is still the top one although all methods improve their task performance. It surpasses the best competing method’s accuracy by a significant margin of 3.3% to 10.0% across all tasks, except for the task of identifying Time of Day (Fig. 6(a)). Effectively, CAL brings the classification model’s average accuracy up to 86.2% after just one iteration of adaptation. With another iteration, CAL boosts the average accuracy to 90.1%, with task accuracy ranging from 83.6% to 96.3%. Through additional iterations, CAL further enhances the accuracy for identifying Weather and Road Function, thus increasing the average accuracy to 92.7%. Figure 6(h) reveals CAL’s superior performance and cost-effectiveness. It consistently outperforms the best competing method in every iteration, with margins ranging from 5.0% to 12.9%. Importantly, CAL has achieved its major improvements in just two iterations, while all other methods require significantly more iterations and so more training data for adaptation.
@@
The importance of CAL’s sample recommendation method is demonstrated by a comparison between Ran-D and CAL, which differ in their sampling strategies. Ran-D uses a simple stratified random sampling strategy, whereas CAL samples data based on features consistency measurement between unlabeled data and the test set for adaptation. CAL consistently improves task accuracy over iterations, whereas Ran-D lowers the accuracy in identifying Road Function and Grade-separated Infrastructure. Moreover, CAL achieves a higher saturated accuracy than Ran-D on all tasks. Consequently, CAL dominates Ran-D in Fig. 6(h).
@@
The groundwork that the driving scene foundation model has laid for multi-attribute scene identification is highlighted by reviewing Ran-I, LLAL-I, and Core-Set-I that do not utilize the foundation model. Those methods need significantly more iterations to improve task performance. Among those, Core-Set-I is the most competitive. After five iterations, it has attained the task accuracy comparable to CAL’s in identifying Weather-related Road Condition and Grade-separated Infrastructure. But CAL still beat Core-Set-I on the remaining five tasks by a margin ranging from 1.3% to 13.9%. Consequently, Core-Set-I’s average accuracy is still 11.8% lower than that of CAL after three iterations and 5.0% lower after five iterations.
@@
[…]
@@
The comparison in Fig. 6 indicates that CAL benefits from the foundation model, which not only sets up a higher starting point for adaptation but also gives CAL good knowledge to identify the most useful training data for ongoing learning and refinement. This pairing of a capable active learning method with a knowledgeable foundation model is the underlying reason for the efficiency and effectiveness of model adaptation.
@@ VI-E Ablation Study
KAA and CAL have now been verified as advantageous designs: KAA builds a knowledge foundation for driving scene identification by acquiring and accumulating information from heterogeneous single-label source data via monotask learning. CAL adapts this foundation model to the multi-label target data. An ablation study was conducted to gauge their respective contributions, with results summarized in TABLE III.
The baseline model in this ablation study is a multitask classification model whose deep encoder is Swin-B pretrained on ImageNet. This baseline model achieves task accuracies ranging from 17.1% to 64.4%, resulting in an average accuracy of 36.6%. Its low accuracy indicates that ImageNet possesses insufficient knowledge regarding driving scene identification.
Adoption of the foundation model as the deep encoder boosts the average accuracy to 67.9%, representing a substantial increase of 31.3%. This gain is attributed to the comprehensive knowledge regarding driving scenes acquired by KAA from diverse single-label datasets. For instance, the accuracy in identifying Road Function, a unique attribute specific to driving scenes, had an increase of 60.3%. Furthermore, the task accuracies for identifying the environmental attributes, Time of Day and Weather, each rise by over 50%. Consequently, the task accuracies now span from 48.1% to 92.8%. However, room for improvement are clearly present. The gains in identifying Weather-related Road Condition and Intersection are limited, primarily due to a large shift in driving scene data distribution compared to the marginal distributions in the single-labeled datasets.
CAL, designed to address the limitation of KAA’s monotask learning approach, effectively improves the average accuracy to 90.1% with only two iterations, corresponding to another increase of 22.2%. At the task level, accuracies have reached the range from 83.6% to 96.3%. With three additional iterations, the task accuracies in identifying Weather and Road Function are improved by at least 5%, further boosting the average accuracy to 92.7%.